my alegebra teacher is wanting me to set up a two variable formula for this word problem help i worked it but i am not sure if it is right i got x is 9 and y is 4 i was hoping you could set it up and let me check mine to see if it is right this is the problem.
The sum of the digits of a two-digit number is 13. If the tens' digit exceeds the units' digit by 5, what is the number?
Thanks for any help you can provide
james

Answer by stanbon(75887)

Let the two digit number be "ab".
a+b= 13 and a=b+5
=Using substitution rewrite the 1st equation as follows:
b+5+b=13
2b+5=13
2b = 8
b = 4 (the unit's digit)
a = b+5 = 4+5=9 (the ten's digit)
You have it right.
Cheers,
stan H.